OpenStudy (anonymous):

factor. 6x^2-13x+6 explain the steps please?

OpenStudy (anonymous):

where did you get that expression?...

OpenStudy (anonymous):

...wat?

OpenStudy (ajprincess):

Let ax^2+bx+c be an expression. a*c=d U need to find two factors, e and f of d such that e+f=-13.K Nw let us come to ur example. a=6 c=6 d=a*c =6*6 =36 factors of 36=(1,2,3,4,6,9,12,18,36,-1,-2,-3,-4,-6,-9,-12,-36) e+f=-13 -4-9=-13 U can split the expression in this way. 6x^2-4x-9x+6 =2x(3x-2)-3(3x-2) =(2x-3)(3x-2) Is that clear?

OpenStudy (maheshmeghwal9):

oops sorry i didn't noticed that; I take back my words.:) Today i m feeling giddy. Deserve a medal @ajprincess

OpenStudy (ajprincess):

It's alright. that sometimes happen. thanx for the medal @maheshmeghwal9.:)

OpenStudy (anonymous):

haha you guys are funny. I need some time to process this... hm LOL i'm stupid so haah

OpenStudy (maheshmeghwal9):

yw & thanx:)

OpenStudy (ajprincess):

yw too.:)

OpenStudy (maheshmeghwal9):

^_^

OpenStudy (maheshmeghwal9):

\[\color{red}{6x^2-13x+6.}\]\[\implies 6x^2-9x-4x+6.\]\[\implies 3x(2x-3)-2(2x-3).\]\[\implies \color{blue}{(3x-2)(2x-3).}\]

OpenStudy (maheshmeghwal9):

so i gt a medal wow. thanx @ajprincess thanx @Steeben for asking such a good question.

OpenStudy (ajprincess):

yw and u deserv it.:)

OpenStudy (ajprincess):

Is that clear @steeben?